One type of sunburn occurs on exposure to UV light of wavelength in the vicinity of 320 nm .

They gave me three questions to this and please help me.
The first is.

What is the energy of one photon of this wavelength in J ?

The second is.

What is the energy of one mole of these photons in J ?

The third is.

How many photons are in a 1.00 mJ burst of this radiation in phtons ?

Please tell me what the heck is all this????
Signed desperate!
Answer #1

1) E =hc/λ

=  (6.626*10^-34*3*10^8)/(320*10^-9)

= 6.21*10^-19 J

2) Of 1 mole E = Avagadro no.*Energy of one photon

= 6.21*10^-19*6.023*10^23

= 374028.3 J

3) n*(6.21*10^-19) = (10^-3)

Thus n = 1.61*10^15